OpenStudy (anonymous):

(y^4)^1/6 Simplify the expression. Assume all variables are positive

OpenStudy (anonymous):

y4/6

OpenStudy (anonymous):

y^2/3

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

4^1/6 = \[\sqrt[6]{4}\] \[y ^{\sqrt[6]{4}}\]

OpenStudy (anonymous):

all you guys are wrong